Calculate the required resistance of an immersion heater that will increase the
temperature of 1.50 kg of water from 10.00C to 50.00C in 10 minutes while operating at
110V.
1
Expert's answer
2021-09-12T19:20:15-0400

"Power = \frac{V^2}{R}"


"= \frac{110^2}{R}"


"= \frac{1.21×10^4}{R}"

"Energy = power \u00d7 time"


"time = 10 minutes = 600 seconds = 6\u00d710^2 s""Energy in Joules =\frac{ (1.21×6)×10^6}{R}
=\frac{7.26 ×10^6}{R} in Joules"


Joules into water = 1.5 kg×4190 J/kg deg×(50-10) deg

= 251,400 Joules =

"=2.51\u00d710^5"

"\frac{7.26 ×10^6 }{R }= 2.51 ×10^5"

"R = \frac{7.26 × 10^6}{2.51× 10^5}"

R = 28.92 ohms